# Long-time Freeness in the Kicked Top

### v1

@@ III Free cumulants dynamics
To study the dynamics at stroboscopic times, we perform an exact diagonalization of the Floquet operator given in Eq. (3), with the parameter fixed at $p=4\pi/7$. Since the Hamiltonian in Eq. (2) is invariant under $\pi$-rotations around the $y$-axis, we restrict our analysis to the positive-parity subspace, which has a dimension of $D=N/2+1$. For concreteness, we focus on the dynamics of the observable
@@
We begin by discussing long-time freeness in the regular, mixed and chaotic regimes (Sec. III.1). Secondly, we focus on the chaotic limit and we show that the dynamics at long times of $2n$-OTOCs of traceless observables is given only by the highest-order free cumulant $\kappa_{2n}(t,0,\dots,t,0)$ (Sec. III.2). Finally, we confirm the validity of ETH in the chaotic regime, showing explicitly the decomposition of multi-time correlation functions in ETH-free cumulants; moreover, we pinpoint the inadequacy of the ETH prediction when it comes to describing the early-time behavior of the square-commutator (Sec. III.3).
III.1 Long-time freeness and chaos
We find that long-time freeness is achieved only in the regime associated with full chaos in the classical limit. This is evident already at the level of the fourth-free cumulant $\kappa_{4}(t,0,t,0)$, expected to vanish in the presence of asymptotic freeness for large $D$, as ${\kappa_{4}(t,0,t,0)\sim\mathcal{O}(D^{-1})}$. In Fig. 2, we report the numerical evaluation of the free cumulant, calculated using Eq. (15c), for different system sizes across three regimes: regular (upper panel, $\gamma=1$), mixed (center panel, $\gamma=4$), and fully chaotic (bottom panel, $\gamma=6$). The data show that in the regular and mixed phase-space regimes, the saturation value remains constant with increasing system size, indicating the absence of long-time freeness. In contrast, for $\gamma=6$, at long times the free cumulant fluctuates around a value that decreases as the Hilbert space dimensionality $D$ increases, as predicted by Eq. (12) for $\hat{B}=\hat{A}(0)$. The absence of late-time freeness in the regular and mixed phase-space can be attributed to the failure of the Eigenstate Thermalization Hypothesis, which is instead satisfied in the chaotic regime, as we discuss in more detail in Sec. III.3.
To quantitatively characterize the emergence of freeness in the chaotic regime, we focus on $\gamma=6$, and we analyze the long-time saturation value of the free cumulants $\kappa_{2n}(t,0\dots,t,0)$ as $n$ and the system size $D$ are changed. We compute the exact time evolution, using Eq. (15), up to $n=6$. In Fig. 3 we consider the time-average $\overline{\kappa_{2n}}$, obtained averaging $\kappa_{2n}(t,0,\dots,t,0)$ over time, between $t=25$ and $t=50$. Free cumulants saturate to a value that vanishes as $D^{-1}$, as encoded in Eq. (12). This is a known feature characterizing long-time dynamics of out-of-time ordered correlators Huang et al. (2019). In the case of free cumulants as well, this value stems from the fluctuations of matrix elements, and as such, it is not presently captured by the full ETH in Eqs. (4). The data show that the plateau is also decreasing with the order of the power $n$, proportionally to $e^{-2n}$. Overall, we find
@@ III.1 Long-time freeness and chaos
After the saturation to $1/D$, one may eventually expect signatures of level repulsion at even longer time scales, see e.g. Ref. Cotler et al. (2017a); Cotler and Hunter-Jones (2020). A more detailed understanding of this phenomenon is left to future investigation.
@@ III.2 Free cumulants and
Free cumulants play a crucial role in the dynamics of the $2n$-out-of-time order correlation $\langle(\hat{A}(t)\hat{A})^{n}\rangle$, as shown in Refs. Pappalardi et al. (2023); Fritzsch and Prosen (2021) for the standard OTOC at $n=2$. We will now discuss it for an arbitrary order of $n$. In particular, we focus on the case $n=4$, but the same results have been obtained for different values of $n$. We consider $2n$-OTOCs of traceless observables, as in Eq. (14), computed via Eq. (15). We will restore the effect of a finite first moment in Sec. III.3 below. Having traceless operators $\langle\hat{A_{0}}\rangle=0$ implies that the $2n$-OTOCs should asymptotically vanish to zero over a long-time since the time-independent plateau has been shifted to zero, see the right-end side of Eq. (13).
@@
$+8\kappa_{6}(t,0,t,0,t,0)\kappa_{2}(t,0)+\kappa_{5}(t,0,t,0,t)\kappa_{3}(t,0,t)+$
$+28\kappa^{2}_{2}(t,0)\kappa_{4}(t,0,t,0)+28\kappa_{2}(t,0)\kappa^{2}_{3}(t,0,t)+$
@@
indicate how quickly freeness is achieved and can be referred as definition of a freeness time-scale. We note that standard $4$-OTOC was already found to decay exponentially fast after the Ehrenfest time in quantum systems with a few-body chaotic limits, see e.g. Refs.García-Mata et al. (2018); Notenson et al. (2023). The goal of Sec. IV will be to interpret the decay rate for arbitrary $n$ in a general way, from a large deviation theory.
III.3 ETH-free cumulants in the chaotic regime
@@
$\langle\hat{A}(t)\hat{A}\hat{A}(t)\hat{A}\rangle_{\mathrm{ETH}}$ $=\mathcal{K}_{4}(t,0,t,0)+\mathcal{K}_{1}^{4}+4\mathcal{K}_{1}^{2}\mathcal{K}_{2}(t,0)+$
$+2\mathcal{K}_{1}^{2}\mathcal{K}_{2}(0,0)+4\mathcal{K}_{1}\mathcal{K}_{3}(0,t,0)+$
$+2\mathcal{K}^{2}_{2}(t,0)\ ,$ (22)
@@
which contains, in turn, out-of-time order correlators. In systems with a well-defined classical chaotic limit, this object grows as Larkin and Ovchinnikov (1969); Kitaev (20145)
@@
in the limit $h_{\rm eff}\to 0$. Beyond this time, quantum effects become significant, causing the square-commutator to saturate. The rate $\tilde{\lambda}$ is often referred to as the quantum Lyapunov exponent García-Mata et al. (2023); Maldacena et al. (2016); Hosur et al. (2016); Xu and Swingle (2022).
@@
In Fig. 6, we compare the numerical results of the exact square-commutator with its ETH-free cumulants decomposition from Eq. (26), which shows very good agreement. By examining the individual free cumulants, a clear hierarchy emerges in their contributions to the square-commutator dynamics: i) $\mathcal{K}^{2}_{2}(0,0)$ determines the saturation value, being the only time-independent free cumulant, which serves as yet another manifestation of freeness; ii) $\mathcal{K}^{2}_{2}(t,0)$ and $\mathcal{K}_{4}(t,t,0,0)$ decay rapidly and fluctuate around zero; iii) $\mathcal{K}_{4}(t,0,t,0)$ drives much of the distinguishing dynamics of the square-commutator, particularly the evolution towards the plateau. Since this term encodes correlations among different matrix elements, the results confirm the idea of ascribing the peculiar dynamics of the square-commutator to the ETH correlations (see, e.g., Ref. Foini and Kurchan (2019a)).
@@
IV Time scale
for the onset of freeness
In Section III.2 we demonstrated that both $2n$-OTOCs and $2n$-free cumulants exhibit exponential decay in reaching long-time freeness, in the kicked-top model, meaning
@@ IV Time scale
The quantities $\tau_{n}$ quantify how quickly freeness is achieved. Building on these findings, we develop a large deviation framework (Sec. IV.1) to understand this time scale, followed by a numerical analysis (Sec. IV.2) in the context of the kicked-top model.
IV.1 Large Deviation Theory for Freeness
@@
where for illustration purposes, we used $R_{n}=n/\tau_{n}$. The $2n$-OTOC are moments of order $n$ of the spectrum of the operator
$\begin{split}\hat{A}_{0}(t)\hat{A}_{0}=\sum_{i}g_{i}(t)|i_{t}\rangle\langle i_{t}|\ ,\\ \text{with}\quad g_{i}(t)=G_{i}e^{-\rho_{i}(t)t}e^{i\theta_{i}(t)}\end{split}$ (29)
where $g_{i}(t)$ are complex eigenvalues, with $e^{-\rho_{i}(t)t}$ representing the time-dependent part of the modulus $|g_{i}(t)|$. In a generic system, the decay rates $\rho_{i}(t)$ are distributed and fluctuate over time. Therefore, probing their higher moments through $2n$-OTOCs provides insight into their dynamical distribution.
To describe their temporal fluctuations, we take inspiration from studies in multifractality, turbulence, and chaos Benzi et al. (1984); Paladin and Vulpiani (1987, 1986); Crisanti et al. (1988); Evers and Mirlin (2000); Macé et al. (2019); Sierant and Turkeshi (2022). After defining the probability distribution of the decay rates as $P(\rho_{t})=\frac{1}{D}\sum^{D}_{i}\delta\left(\rho_{t}-\rho_{i}(t)\right)$, we assume the existence of a typical value $\rho_{\rm typ}$ and a Large Deviation form at long-times for it:
@@
$\langle(\hat{A}_{0}(t)\hat{A}_{0})^{n}\rangle$ $=\frac{1}{D}\sum_{i}G^{n}_{i}e^{-n\rho_{i}(t)t}e^{in\theta_{i}}$
@@
IV.2 Freeness multifractality in the Kicked Top
@@
As discussed in Sec. III.1 above, at longer times, the $2n$-OTOCs display fluctuations around a plateau $\mathcal{O}(D^{-1})$ indicating freeness.
@@
This is shown in Fig. 9, where we plot the time $t_{\epsilon}$ at which each $2n$-OTOC reaches an $\epsilon$ value, as a way to quantify the shift. For definiteness, we fix $\epsilon=-2.55$ and calculate $t_{\epsilon}$ as the first interpolated time for which $\log\left|\langle(\hat{A}_{0}(t)\hat{A}_{0})^{n}\rangle\right|/n\leq\epsilon$, for $n=1,2,3,4$. The data show a clear logarithmic increase of $t_{\epsilon}$ for even $n$, while the time is constant for odd $n$. Furthermore, we also note that the time shift depends on the amount of chaos present in the system. We calculate the time shift also for $\gamma=4$ and $n=2$ corresponding to the mixed phase-space region discussed, see e.g. Fig. 2 in Sec. III.1. In this case, the time shift $t_{\epsilon}$ is bigger than the corresponding one for $\gamma=6$, indicating a smaller shift for more chaotic systems.
In the case of models with a semiclassical chaotic limit $\hbar_{\rm eff}\to 0$, the explanation for this effect can be found in the early-time dynamics of the $2n$-OTOCs. This behavior has been studied for few-body quantum chaos García-Mata et al. (2018); Notenson et al. (2023), as well as for the Sachdev-Ye-Kitaev (SYK) model Kobrin et al. (2021), for $n=2$. In this case, the OTOC is known to have a Lyapunov regime at short intermediate times, meaning it is governed by the exponential growth of the square-commutator (cf. Eq. (24)). In particular,
@@
For higher and even $n=2m$, a similar behavior to the one of Eq. (35a) is conjectured to be valid at short intermediate times Pappalardi and Kurchan (2023), i.e.
@@
for $1\ll t\ll t^{(2m)}_{\rm Ehr}$, where $L_{2m}$ are the rates accounted by the so-called generalized Lyapunov exponents, which stem from the powers of the square-commutator ${\langle|[\hat{A}(t),\hat{A}(0)]|^{2m}\rangle\sim\hbar_{\rm eff}^{2m}e^{L_{2m}t}}$. One retrieves the square commutator for $m=1$ ($n=2$). This behavior is valid until the generalized Ehrenfest time ${t^{(2m)}_{\rm Ehr}\sim\log\hbar_{\rm eff}^{-2m}/L_{2m}}$, which is assumed to be increasing with $m$, see also Refs.Trunin (2023a, b). After this time scale, the higher powers of the square-commutator saturate. In the case of odd powers, this effect is not present, indeed, the odd powers of the commutator usually do not have exponential growth due to a cancellation of signs.

### v2

@@ 3 Free cumulants dynamics
To study the dynamics at stroboscopic times, we perform an exact diagonalization of the Floquet operator given in Eq. (3), with the parameter fixed at $p=4\pi/7$. Since the Hamiltonian in Eq. (2) is invariant under $\pi$-rotations around the $y$-axis, we restrict our analysis to the positive-parity subspace, which has a dimension of $D=N/2+1$ when $j=N/2$ is even [^6] When $j=N/2$ is odd the dimension is $D=N/2$, while if $j=N/2$ is half-integer the dimension is $D=\frac{1}{2}(N+1)$. . For concreteness, we focus on the dynamics of the observable
@@
We begin by discussing long-time freeness in the regular, mixed and chaotic regimes (Sec. 3.1). Secondly, we focus on the chaotic limit and we show that the dynamics at long times of $2n$-OTOCs of traceless observables is given only by the highest-order free cumulant $\kappa_{2n}(t,0,\dots,t,0)$ (Sec. 3.2). Finally, we confirm the validity of ETH in the chaotic regime, showing explicitly the decomposition of multi-time correlation functions in ETH-free cumulants; moreover, we pinpoint the inadequacy of the ETH prediction when it comes to describing the early-time behavior of the square-commutator (Sec. 3.3).
3.1 Long-time freeness and chaos
We find that long-time freeness is achieved only in the regime associated with full chaos in the classical limit. This is evident already at the level of the fourth-free cumulant $\kappa_{4}(t,0,t,0)$, expected to vanish in the presence of asymptotic freeness for large $D$, as ${\kappa_{4}(t,0,t,0)\sim\mathcal{O}(D^{-1})}$. In Fig. 2, we report the numerical evaluation of the free cumulant, calculated using Eq. (15c), for different system sizes across three regimes: regular (upper panel, $\gamma=1$), mixed (center panel, $\gamma=4$), and fully chaotic (bottom panel, $\gamma=6$). The data show that in the regular and mixed phase-space regimes, the saturation value remains constant with increasing system size, indicating the absence of long-time freeness. In contrast, for $\gamma=6$, at long times the free cumulant fluctuates around a value that decreases as the Hilbert space dimensionality $D$ increases, as predicted by Eq. (12) for $\hat{B}=\hat{A}(0)$. The absence of late-time freeness in the regular and mixed phase-space can be attributed to the failure of the Eigenstate Thermalization Hypothesis, which is instead satisfied in the chaotic regime, as we discuss in more detail in Sec. 3.3.
To quantitatively characterize the emergence of freeness in the chaotic regime, we focus on $\gamma=6$, and we analyze the long-time saturation value of the free cumulants $\kappa_{2n}(t,0\dots,t,0)$ as $n$ and the system size $D$ are changed. We compute the exact time evolution, using Eq. (15), up to $n=6$. In Fig. 3 we consider the time-average $\overline{\kappa_{2n}}$, obtained averaging $\kappa_{2n}(t,0,\dots,t,0)$ over time, between $t=25$ and $t=50$. Free cumulants saturate to a value that vanishes as $D^{-1}$, as encoded in Eq. (12). This is a known feature characterizing long-time dynamics of out-of-time ordered correlators [107]. In the case of free cumulants as well, this value stems from the fluctuations of matrix elements, and as such, it is not presently captured by the full ETH in Eqs. (4). The data show that the plateau is also decreasing with the order of the power $n$, proportionally to $e^{-2n}$. Overall, we find
@@ 3.1 Long-time freeness and chaos
After the saturation to $1/D$, one may eventually expect signatures of level repulsion at even longer time scales, see e.g. Ref. [33, 35]. A more detailed understanding of this phenomenon is left to future investigation.
@@ 3.2 Free cumulants and
Free cumulants play a crucial role in the dynamics of the $2n$-out-of-time order correlation $\langle(\hat{A}(t)\hat{A})^{n}\rangle$, as shown in Refs. [58, 108] for the standard OTOC at $n=2$. We will now discuss it for an arbitrary order of $n$. In particular, we focus on the case $n=4$, but the same results have been obtained for different values of $n$. We consider $2n$-OTOCs of traceless observables, as in Eq. (14), computed via Eq. (15). We will restore the effect of a finite first moment in Sec. 3.3 below. Having traceless operators $\langle\hat{A_{0}}\rangle=0$ implies that the $2n$-OTOCs should asymptotically vanish to zero over a long-time since the time-independent plateau has been shifted to zero, see the right-end side of Eq. (13).
@@
$+8\kappa_{6}(t,0,t,0,t,0)\kappa_{2}(t,0)+$
$+\kappa_{5}(t,0,t,0,t)\kappa_{3}(t,0,t)+$
$+28\kappa^{2}_{2}(t,0)\kappa_{4}(t,0,t,0)+$
$+28\kappa_{2}(t,0)\kappa^{2}_{3}(t,0,t)+$
@@
indicate how quickly freeness is achieved and can be referred as definition of a freeness time-scale. We note that standard $4$-OTOC was already found to decay exponentially fast after the Ehrenfest time in quantum systems with a few-body chaotic limits, see e.g. Refs.[109, 110]. The goal of Sec. 4 will be to interpret the decay rate for arbitrary $n$ in a general way, from a large deviation theory.
3.3 ETH-free cumulants in the chaotic regime
@@
$\langle\hat{A}(t)\hat{A}\hat{A}(t)\hat{A}\rangle_{\mathrm{ETH}}$ $=\mathcal{K}_{4}(t,0,t,0)+\mathcal{K}_{1}^{4}+$ (22)
$+4\mathcal{K}_{1}^{2}\mathcal{K}_{2}(t,0)+2\mathcal{K}_{1}^{2}\mathcal{K}_{2}(0,0)+$
$+4\mathcal{K}_{1}\mathcal{K}_{3}(0,t,0)+2\mathcal{K}^{2}_{2}(t,0)\ ,$
@@
which contains, in turn, out-of-time order correlators. In systems with a well-defined classical chaotic limit, this object grows as [30, 31]
@@
in the limit $h_{\rm eff}\to 0$. Beyond this time, quantum effects become significant, causing the square-commutator to saturate. The rate $\tilde{\lambda}$ is often referred to as the quantum Lyapunov exponent [29, 26, 27, 28].
@@
In Fig. 6, we compare the numerical results of the exact square-commutator with its ETH-free cumulants decomposition from Eq. (26), which shows very good agreement. By examining the individual free cumulants, a clear hierarchy emerges in their contributions to the square-commutator dynamics: i) $\mathcal{K}^{2}_{2}(0,0)$ determines the saturation value, being the only time-independent free cumulant, which serves as yet another manifestation of freeness; ii) $\mathcal{K}^{2}_{2}(t,0)$ and $\mathcal{K}_{4}(t,t,0,0)$ decay rapidly and fluctuate around zero; iii) $\mathcal{K}_{4}(t,0,t,0)$ drives much of the distinguishing dynamics of the square-commutator, particularly the evolution towards the plateau. Since this term encodes correlations among different matrix elements, the results confirm the idea of ascribing the peculiar dynamics of the square-commutator to the ETH correlations (see, e.g., Ref. [43]).
@@
4 Time scale for the onset of freeness
In Section 3.2 we demonstrated that both $2n$-OTOCs and $2n$-free cumulants exhibit exponential decay in reaching long-time freeness, in the kicked-top model, meaning
@@ 4 Time scale for the onset of freeness
The quantities $\tau_{n}$ quantify how quickly freeness is achieved. Building on these findings, we develop a large deviation framework (Sec. 4.1) to understand this time scale, followed by a numerical analysis (Sec. 4.2) in the context of the kicked-top model.
4.1 Large Deviation Theory for Freeness
@@
where for illustration purposes, we used $R_{n}=n/\tau_{n}$. The $2n$-OTOC can be seen as moments of order $n$ of the spectrum of the operator $\hat{A}_{0}(t)\hat{A}_{0}$: we call $g_{i}(t)$ its complex eigenvalues [^8] Note that the eigenvalues of $\hat{A}_{0}(t)\hat{A}_{0}$ come in conjugate pairs, ensuring that the trace of its moments is a real number. This property follows from the fact that $\hat{A}_{0}(t)\hat{A}_{0}$ is a pseudo-hermitian operator, i.e. it obeys $(A_{0}(t)A_{0})^{\dagger}=\eta(A_{0}(t)A_{0})\eta^{-1}$ with $\eta=A_{0}$. and we assume that the modulus $|{g_{i}(t)}|$ has the time-dependent part given by $e^{-\rho_{i}(t)t}$. Specifically:
$g_{i}(t)=G_{i}e^{-\rho_{i}(t)t}e^{i\theta_{i}(t)}\ .$ (29)
In a generic system, the decay rates $\rho_{i}(t)$ are distributed and fluctuate over time. Therefore, probing their higher moments through $2n$-OTOCs provides insight into their dynamical distribution.
To describe their temporal fluctuations, we take inspiration from studies in multifractality, turbulence, and chaos [88, 89, 90, 91, 92, 93, 94]. After defining the probability distribution of the decay rates as $P(\rho_{t})=\frac{1}{D}\sum^{D}_{i}\delta\left(\rho_{t}-\rho_{i}(t)\right)$, we assume the existence of a typical value $\rho_{\rm typ}$ and a Large Deviation form at long-times for it:
@@
$\langle(\hat{A}_{0}(t)\hat{A}_{0})^{n}\rangle$ $=\frac{1}{D}\sum_{i}g^{n}_{i}=\frac{1}{D}\sum_{i}G^{n}_{i}e^{-n\rho_{i}(t)t}e^{in\theta_{i}(t)}$
@@
4.2 Freeness multifractality in the Kicked Top
@@
As discussed in Sec. 3.1 above, at longer times, the $2n$-OTOCs display fluctuations around a plateau $\mathcal{O}(D^{-1})$ indicating freeness.
@@
This is shown in Fig. 9, where we plot the time $t_{\epsilon}$ at which each $2n$-OTOC reaches an $\epsilon$ value, as a way to quantify the shift. For definiteness, we fix $\epsilon=-2.55$ and calculate $t_{\epsilon}$ as the first interpolated time for which $\log\left|\langle(\hat{A}_{0}(t)\hat{A}_{0})^{n}\rangle\right|/n\leq\epsilon$, for $n=1,2,3,4$. The data show a clear logarithmic increase of $t_{\epsilon}$ for even $n$, while the time is constant for odd $n$. Furthermore, we also note that the time shift depends on the amount of chaos present in the system. We calculate the time shift also for $\gamma=4$ and $n=2$ corresponding to the mixed phase-space region discussed, see e.g. Fig. 2 in Sec. 3.1. In this case, the time shift $t_{\epsilon}$ is bigger than the corresponding one for $\gamma=6$, indicating a smaller shift for more chaotic systems.
In the case of models with a semiclassical chaotic limit $\hbar_{\rm eff}\to 0$, the explanation for this effect can be found in the early-time dynamics of the $2n$-OTOCs. This behavior has been studied for few-body quantum chaos [109, 110], as well as for the Sachdev-Ye-Kitaev (SYK) model [111], for $n=2$. In this case, the OTOC is known to have a Lyapunov regime at short intermediate times, meaning it is governed by the exponential growth of the square-commutator (cf. Eq. (24)). In particular,
@@
For higher and even $n=2m$, a similar behavior to the one of Eq. (35a) is conjectured to be valid at short intermediate times [37], i.e.
@@
for $1\ll t\ll t^{(2m)}_{\rm Ehr}$, where $L_{2m}$ are the rates accounted by the so-called generalized Lyapunov exponents, which stem from the powers of the square-commutator ${\langle|[\hat{A}(t),\hat{A}(0)]|^{2m}\rangle\sim\hbar_{\rm eff}^{2m}e^{L_{2m}t}}$. One retrieves the square commutator for $m=1$ ($n=2$). This behavior is valid until the generalized Ehrenfest time ${t^{(2m)}_{\rm Ehr}\sim\log\hbar_{\rm eff}^{-2m}/L_{2m}}$, which is assumed to be increasing with $m$, see also Refs.[38, 39]. After this time scale, the higher powers of the square-commutator saturate. In the case of odd powers, this effect is not present, indeed, the odd powers of the commutator usually do not have exponential growth due to a cancellation of signs.